OpenStudy (anonymous):

ill fan and give medals Evaluate the function f(x) = 2x + 2, when x = –1.

OpenStudy (anonymous):

2*-1+2=0

OpenStudy (anonymous):

it equals 0

OpenStudy (anonymous):

@nater can you help me with one more ?

OpenStudy (the_fizicx99):

\[t = 50\] \[g(t) = \frac{ 4 }{ 5}t\] Plug in 50 for t \[g(50) = \frac{ 4 }{ 5 } \times \frac{ 50 }{ 1 }\] \[g(50) = \frac{ 4 }{ 5}\times \frac{ 50 }{ 1 } \rightarrow \frac{ 200 }{ 5 } \rightarrow 40\]